Grade 11General Physics

A body of mass 1 kg is pushed up an incline surface;inclined at 30 degree to the horizontal,by a force of 10N parallel to the inclined surface.If-the coefficient of friction between block and the incline is 0.1 and the block is pushed up by 10m along the incline;calculate th work done against he gravity.

Vidit
10 Years agoGrade 11

1 Answer

Abhishek Singh
ApprovedApproved Tutor Answer10 Years ago
the work done agains the gravity is = mg(h*sin(theta)) = 1*10*10*1/2 = 50 joules.
 
You should not consider the other force as you have to calculate the work done aginst gravity only, so focus on gravity only